OpenStudy (anonymous):

@campbell_st Cards numbered 1, 1, 2, 3, 4, 4, 4, 5, 6, 6, 7, 8, 8, 8, and 9 are placed in a bag. If a card is drawn from the bag and then replaced 135 times, about how many times would a 6 be drawn? A. 2 B. 9 C. 15 D. 18

OpenStudy (anonymous):

Help please i think it is D am i right

OpenStudy (campbell_st):

well its the probability of drawing 1 six.... times the number of draws... P(6) x 135 = hope it helps so P(6) = 2/15 and D is a great answer

OpenStudy (anonymous):

can u help with more please

OpenStudy (anonymous):

Valerie plays a game using the spinner shown. During the game she will spin a total of 36 times. Which is the best prediction for the number of times she will get either "lose a turn" or "back 3" on the spinner? A. 2 B. 6 C. 12 D. 18

OpenStudy (anonymous):

@campbell_st help

OpenStudy (anonymous):

i think it is B. am i right

OpenStudy (campbell_st):

well there is 1 lose a turn and 1 back 3 so P(lose or back 3) = 1/3 so in 36 spins expected No = 1/3 x 36 hope it helps

OpenStudy (anonymous):

that equals C.

OpenStudy (campbell_st):

thats right, well done